Need help. Use the method of substitution to solve the system of linear equations. 2x-y=8 9x-4y=36.
Can you please show work.

Answer by stanbon(75887)

Use the method of substitution to solve the system of linear equations.
2x-y=8
9x-4y=36.
-----
Solve the 1st equation for "y":
y = 2x-8
---
Substitute for "y" and solve for "x":
9x - 4(2x-8) = 36
9x - 8x + 32 = 36
x = 4
----
Solve for "y":
y = 2x-8
y = 2*4-8
y = 0
-------
Cheers,
Stan H.
===========